# MoLA: Motion Generation and Editing with Latent Diffusion Enhanced by Adversarial Training

### v2

@@ Appendix C Evaluation metrics details
$S_{d}=300$ is used in our experiments.
@@
$S_{l}=10$ is used in our experiments.
Appendix D Additional experiments for motion generation
Here, we discuss the effect of adversarial training on Stage 2 to supplement the experiments in Sections 4.3 and 4.4. To demonstrate the broad effectiveness of adversarial training in text-to-motion generation, we apply adversarial training not only to a continuous latent case as in MLD (Chen et al., 2023) introduced in Section 3 but also to a discrete latent case as in T2M-GPT (Zhang et al., 2023b) introduced in Sections 4.2 and 4.3. The results of two types of methods, (top:) a discrete latent case and (bottom:) a continuous latent case, are shown in Tables 7 and 8. As shown in the tables, our proposed strategy improves existing methods in FID, which is a metric that evaluates the overall motion quality, on both continuous latent and discrete latent cases on both HumanML3D (Guo et al., 2022) and KIT-ML (Plappert et al., 2016) datasets. Moreover, we can see replacing GAN with SAN boosts performance more on FID, R-precision, and MMDist metrics on both datasets. Note that the only difference between baseline models and our models is training method, not inference, as mentioned in Section 4.4. Therefore, inference with our model can be achieved with equivalent time to MLD and T2M-GPT baselines.

### v3

@@ Appendix D Evaluation metrics details
$S_{d}=300$ is used in our experiments, following previous works.
@@
$S_{l}=10$ is used in our experiments, following previous works.
Appendix E Additional experiments
E.1. Performance comparison on the KIT-ML dataset
We evaluate our method not only on the HumanML3D dataset presented in Section 4.1 but also on the KIT-ML dataset (Plappert et al., 2016), comparing its performance against other text-to-motion generation methods. The KIT-ML dataset includes 3,911 human motion sequences and 6,278 text descriptions derived from the KIT (Mandery et al., 2015) and CMU (CMU, 2003) datasets. These data are split into training, validation, and test sets with proportions of 80%, 5%, and 15%, respectively. Consistent with Section 4.1, the evaluated methods are categorized into three groups: (i) those using VQ-based latent representations (discrete), (ii) those employing data-space diffusion models (continuous, raw data), and (iii) those utilizing VAE-based latent representations (continuous, latent). As shown in Table 4, MoLA achieves the best performance in terms of R-Precision and MMDist, particularly among continuous-based methods. However, it does not outperform methods like MLD in terms of FID. Comparing this with Table 1, we observe a discrepancy in the FID trends across the two datasets, suggesting there is room for improvement in this aspect of MoLA.
Top-1 Top-2 Top-3
MoMask (Guo et al., 2024) $0.433^{\pm.007}$ $0.656^{\pm.005}$ $0.781^{\pm.005}$ $0.204^{\pm.011}$ $2.779^{\pm.022}$ $1.131^{\pm.043}$
MMM (Pinyoanuntapong et al., 2024b) $0.381^{\pm.005}$ $0.590^{\pm.006}$ $0.718^{\pm.005}$ $0.429^{\pm.019}$ $3.146^{\pm.019}$ $10.633^{\pm.097}$ $1.105^{\pm.026}$
ParCO (Zou et al., 2024) $0.430^{\pm.004}$ $0.649^{\pm.007}$ $0.772^{\pm.008}$ $0.453^{\pm.027}$ $2.820^{\pm.028}$ $10.95^{\pm.094}$ $1.245^{\pm.022}$
BAMM (Pinyoanuntapong et al., 2024a) $0.438^{\pm.009}$ $0.661^{\pm.009}$ $0.788^{\pm.005}$ $0.183^{\pm.013}$ $2.723^{\pm.026}$ $11.008^{\pm.094}$ $1.609^{\pm.065}$
Continuous (latent) MLD (Chen et al., 2023) $0.390^{\pm.008}$ $0.609^{\pm.008}$ $0.734^{\pm.007}$ $\bm{0.404^{\pm.027}}$ $3.204^{\pm.027}$ $10.80^{\pm.117}$ $\bm{2.192^{\pm.071}}$
MotionLCM (Dai et al., 2024) -- -- -- -- -- -- --
MoLA (ours) $\bm{0.432^{\pm.008}}$ $\bm{0.655^{\pm.008}}$ $\bm{0.770^{\pm.004}}$ $0.529^{\pm.056}$ $\bm{2.942^{\pm.053}}$ $11.129^{\pm.158}$ $1.789^{\pm.174}$
Table 4. Comparison with state-of-the-art methods on KIT-ML. Note that discrete representations do not allow for training-free motion editing; therefore, methods based on VQ-based latent representations (Discrete) are grayed out. The best scores for each metric in the methods using VAE-based latent representations (Continuous (latent)) are highlighted in bold.
E.2. Ablation of classifier-free diffusion guidance on stage 2
As explained in Section 3.3, we employ a classifier-free diffusion guidance technique (Ho and Salimans, 2022). In training, we randomly drop the condition $\tau(\bm{c})$ with a probability of 10% and train both the conditional model $\epsilon_{\theta}(z_{t},t,\tau(\bm{c}))$ and the unconditional model $\epsilon_{\theta}(z_{t},t,\varnothing)$. In inference, the two predictions are linearly combined as follows:
(19) $\epsilon_{\theta}(z_{t},t,\bm{c})=s\cdot\epsilon_{\theta}(z_{t},t,\tau(\bm{c}))+(1-s)\cdot\epsilon_{\theta}(z_{t},t,\varnothing),$
where $s$ is the guidance scale, and $s>1$ can amplify the effect of the guidance. We investigate how the performance of our method changes with different values of the hyperparameter $s$. From Figure 8, we observe that the best performance in FID is achieved at $s=11$ on the HumanML3D dataset. Consequently, we adopt $s=11$ in Section 4.1.
Figure 8. Evaluation sweep over guidance scale $s$